OpenStudy (anonymous):

1g of cork with a charge 1mueC floats motionless 1cm above a large uniformly charged plane near the surface of the earth. the surface charge density of the plane assuming it to be infinite sheet is

OpenStudy (experimentx):

F = 1 * 10^-3 * 9.8 = 1 * 10^-6 * E, where E = http://hyperphysics.phy-astr.gsu.edu/hbase/electric/elesht.html